Shift report 20030314 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Mar 14 2003 - 08:02:47 EST

  • Next message: Apache: "Shift report 20030314 08:00-16:00"
    ------------------------------------------------------------
    Date: Fri Mar 14 08:02:47 2003, Shift: 00:00-08:00
    Supervisor: CEJ, with: 
    Log book#14 pages -1 - -1
    
    Shift Summary:
    ------------------------------------------------------------
    1) End of store at around 1:30 (gap cleaning from 00:30!). 
       New store at 03:00. BRAHMS ZDC rates 3/4 of PHOBOS, 1/2 
       of PHENIX. Asked for steering at 05:00 - that didn't help.
    
    2) Ped, TPC ped and DVM runs 8362, 8363 and 8364.
       MRS@90(B1050), FS@4(A1/1)
       8361 Trig 3 107K, trig 2 7.6K (started in prev shift)
       8365 Trig 3 ~200K?, trig 2 18K
       8367 Trig 3 140K, trig 2 10K
       8368 Trig 3  22K, trig 2 1.5K
    
    3) The lum scaler monitor showed zero MRS coins, but the 
       data was there. No detector problems reported.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri Mar 14 2003 - 08:07:30 EST